OpenStudy (anonymous):

Find the x intercept for: y=2-√x / 5x

OpenStudy (astrophysics):

Set y = 0 and solve for x

OpenStudy (astrophysics):

\[y=\frac{ 2-\sqrt{x} }{ 5x }\] is it this

OpenStudy (anonymous):

Yes

OpenStudy (astrophysics):

Ok let y = 0 and solve for x as I suggested

OpenStudy (astrophysics):

\[\frac{ 2-\sqrt{x} }{ 5x }=0\]

OpenStudy (anonymous):

So 5x=2-√x

OpenStudy (astrophysics):

Not quite, remember you have to multiply both sides by 5x, so we have 5x * 0 = 0 right? So \[2-\sqrt{x}=0\]

OpenStudy (anonymous):

Thank you! I just couldn't see what i was missing but its multiplication of both sides not subtraction -.-

OpenStudy (anonymous):

Very appreciated!

OpenStudy (astrophysics):

Np, so lets just finish it off, what did you get

OpenStudy (anonymous):

-√x = 2 square both sides -x = 4 x+-4

OpenStudy (anonymous):

x = -4

OpenStudy (astrophysics):

Well it would be better if you just move \[-\sqrt{x}\] to the right side so we have \[2-\sqrt{x}=0 \implies 2 = \sqrt{x} \implies x = 2^2 = 4\]

OpenStudy (anonymous):

Maybe i'm just not looking at everything correct today =/ I'm sorry!

OpenStudy (astrophysics):

It's cool